OpenStudy (anonymous):

solve for c 1/2(sqrt c-2) = sqrt 5/5

OpenStudy (anonymous):

need algebra help

OpenStudy (anonymous):

\[1/(2(\sqrt{c-2))}=\sqrt{5}/5\]

OpenStudy (whpalmer4):

\[\frac{1}{2\sqrt{c-2}}=\frac{\sqrt{5}}{5}\]Is that correct? Have you tried cross-multiplication and solving for \(c\)?

OpenStudy (lovelyanna):

@whpalmer4 When you finish here, help me with my question?

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

not getting right answer

OpenStudy (anonymous):

cross multiply = \[2(\sqrt{5(c-2)}/5\]

OpenStudy (whpalmer4):

No. \[\frac{a}{b} = \frac{c}{d}\]cross-multiply: \[a*d = b*c\] I don't see an = anywhere in your work

OpenStudy (anonymous):

\[5=2\sqrt{5(c-2)}\] \[\sqrt{5}=2[5(c-2)]\]

OpenStudy (anonymous):

does that look better so far?

OpenStudy (anonymous):

sqrt5=10-2c-4 not sure what to do

OpenStudy (anonymous):

6sgrt5/2=c=3sqrt5

OpenStudy (anonymous):

which is incorrect

OpenStudy (whpalmer4):

\[\frac{1}{2\sqrt{c-2}}=\frac{\sqrt{5}}{5}\]\[5*1 = \sqrt{5}*2\sqrt{c-2}\]\[5 = 2\sqrt{5}\sqrt{c-2}\]

OpenStudy (whpalmer4):

At this point, why don't you square both sides?

OpenStudy (anonymous):

i did I though? sqrt5=2[5(c-2)]

OpenStudy (whpalmer4):

\[5*5 = 2*2*\sqrt{5}*\sqrt{5}*\sqrt{c-2}*\sqrt{c-2}\]\[25 = 4*5*(c-2)\]

OpenStudy (anonymous):

sqrt 5=10-2c-4

OpenStudy (anonymous):

oh

OpenStudy (whpalmer4):

\(a\) squared = \(a^2 = a*a\)

OpenStudy (anonymous):

25=20(c-2) 25=20c-40 65=20c 13/4=c

OpenStudy (whpalmer4):

\[(\sqrt{a})^2 = a\](assuming \(a\ge0\))

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

was t

OpenStudy (anonymous):

was taking sqrt root instead of squaring

OpenStudy (anonymous):

how do I know when to use correct method?

OpenStudy (whpalmer4):

Yes, now you should check your answer in the original equation. Whenever you square both sides to eliminate a radical, you need to test any solutions to make sure they are not extraneous (read: don't work in the original equation).

OpenStudy (anonymous):

original q is find c mean value theorem f(x)=(sqrt x-2) [2,7]

OpenStudy (anonymous):

so 13/4 = 3 1/4 is in interval is that what you mean by checking in original question

OpenStudy (whpalmer4):

No, I mean you plug c = 13/4 into the original equation, and make sure the result is true. It isn't an issue in this problem, because you had a radical on both sides, but if you solve something like \[2\sqrt{x-1} = x-3\] you get two solutions, but only one of them works as a solution to the original equation.

OpenStudy (anonymous):

ok thanks

OpenStudy (lovelyanna):

@whpalmer4 Waiting whenever your ready :-)